# The Properties of Little Red Dot Galaxies in the ASTRID Simulation

### v1

@@ 5 Where Are The Dimmer LRDs?
The population of LRDs in Astrid represents a subset of the observed population, as their number density is significantly lower than those found in observations, and they are concentrated within the brighter portion of the observed LRD population. Astrid contains $\sim 1/100$ of the expected number of LRDs with UV magnitudes brighter than $\rm M_{UV}=-19$, but completely lacks LRDs with $\rm M_{UV}\geq-18.5$ resulting in an overall LRD number density of $\sim 1/1000$ of observations (Kocevski et al., 2024; Kokorev et al., 2024; Pérez-González et al., 2024)
Figure 11: Spectrum of a $M_{*}=10^{9.4}\,M_{\odot}$ galaxy hosting a black hole with mass $M_{BH}=10^{6.1}\,M_{\odot}$, accreting at 116% the Eddington rate, with the dust occlusion on each star increased by a factor of four and applying an AGN SED model with a Hydrogen column density of $N_{H}=10^{24}$ $\rm N/cm^{2}$. This is the same galaxy that appears in the bottom left panel of Figure 6. With these changes, the galaxy fits the criteria to be an LRD.
@@
This galaxy has an F277W-F [^4] W color of 0.88, which is insufficient to host an LRD. The primary differences between the spectrum of this galaxy and that of the example LRD in Figure 2 are the relative brightness of the AGN and the dust attenuation of both components. In the LRD spectrum, before dust attenuation, the AGN and stellar components have similar brightness in the F [^4] W band. Despite having one of the brightest AGN relative to its stellar mass among the lower stellar mass galaxies, the AGN emission of the dim LRD candidate galaxy is only $\sim 16\%$ as bright as its host galaxy in the F [^4] W band before dust attenuation. This indicates very few, if any, of the lower mass galaxies in Astrid have AGN that are bright enough to significantly impact their overall color. This is likely the result of both black hole accretion and star formation being induced by the presence of cold dense gas, producing host galaxies which are brighter due to their increased stellar mass, and young stellar population. This coupling is broken once AGN feedback begins quenching the host galaxy, as we see in our LRDs and models of high redshift galaxies hosting over-massive black holes (Pacucci & Loeb, 2024; Pacucci et al., 2024).
@@
These two factors are heavily impacted by modeling choices in our mock observation pipeline. AGN bolometric luminosity is determined by the simulation, but the AGN SED can impact how that luminosity is distributed, and is dependent on the properties of the gas very near the black hole, which is not resolved by cosmological simulations. Recent work (e.g., Naidu et al., 2025), suggests that AGN embedded in a denser gas cloud, similar to models of direct collapse black holes (e.g. Pacucci et al., 2015, 2016), may be a strong candidate for LRD emission, which produces a significantly different SED than our more traditional model. Similarly, our dust model has a normalization factor calibrated in Bird et al. (2022), but may not be tuned appropriately for LRD galaxies.
@@
We experiment with these two modeling choices on our dim LRD candidate galaxy. We increase the hydrogen column density of our AGN SED model in cloudy from $10^{23}$ $\rm N/cm^{2}$ to $10^{24}$ $\rm N/cm^{2}$ to approximate the effect of the "black hole star" model proposed in Naidu et al. (2025). Additionally, we increase the dust normalization by a factor of four (from $\kappa_{\rm ISM}=10^{4.1}$ to $\kappa_{\rm ISM}=10^{4.7}$) for the stellar emission to explore the impact of errors in normalization. This also serves as an approximate test of the impact of potential under-density due to simulation effects like resolution limitations and gravitational softening.
[…]
@@
The increased hydrogen column density in the AGN SED introduces both a steeper intrinsic AGN SED and a significant Balmer break, both of which contribute to the red appearance of this galaxy. These qualities are especially important for any LRDs with minimal dust content, like those suggested by Setton et al. (2025); Xiao et al. (2025), and would result in the consistent Balmer break transition that is found in Setton et al. (2024).
@@
The increased stellar dust attenuation reddens the stellar contribution to the total spectrum and reduces the amount of stellar emission relative to the AGN emission, allowing for the AGN to have a larger impact on the total spectrum. While increased dust attenuation is one path to a dimmer stellar emission, if a similar black hole were present in a lower mass host galaxy, that would also facilitate this effect. With the black hole models used in Astrid this is unlikely to occur due to AGN feedback, but it may be possible under different black hole models. The BRAHMA simulation suite (Bhowmick et al., 2024) suggests that the over-massive black holes seen in observations are reproducible in simulation with specific seeding techniques, and that much of their mass is the result of black hole mergers. This would allow for the formation of more massive black holes in lower mass galaxies, as their mass growth would be less coupled to star formation. These more massive black holes would be capable of brighter AGN emission without exceeding the Eddington limit, potentially appearing as LRDs.
@@ 6.1 Discussion
The size of the Astrid simulation is exceptionally advantageous for works such as this one, as it provides a cosmological volume that is much larger than most observational datasets. Given the density of LRDs found by observations across the redshifts we analyze in this work, we would expect to identify tens of thousands of LRDs in our dataset (Kocevski et al., 2024; Kokorev et al., 2024; Matthee et al., 2023; Pérez-González et al., 2024). Despite this, we find only 17 LRDs, and 24 no-break LRDs in our dataset, and those galaxies would be among the brightest LRDs observed.
Due to the compact nature of LRDs, the computational constraints that arise on small scales in cosmological simulations may limit their ability to form. The SPH method used in Astrid allows for highly accurate physical modeling on a wide range of scales. Still, it is limited at the smallest scales, due to the mass resolution of particles and gravitational softening. This is relevant for all LRD models, but is of particular importance to the star-dominated models because they are reliant on the existence of extremely dense star-forming regions, which can push the theoretical limits of the current cosmological model (Guia et al., 2024; Ma et al., 2024; Leung et al., 2024; Akins et al., 2024).
Simulations also must choose specific physical models for processes regarding black holes (e.g., seeding, accretion, and merging) and gas dynamics/radiative properties (e.g., AGN feedback, star formation), etc., all of which impact the possibility of an LRD forming. The choices for these models in Astrid are described in detail in Ni et al. (2022); Bird et al. (2022) and have been shown to have results consistent with observations and other simulations on a wide variety of metrics. AGN-dominated fits of observed LRDs often suggest that over-massive black holes must be present to produce AGNs with sufficient brightness (Pacucci & Loeb, 2024; Pacucci et al., 2023; Inayoshi & Ichikawa, 2024; Durodola et al., 2025), but most current black hole seeding and accretion models, including those used in Astrid  do not have the pathways to create such over-massive black holes.
@@
Despite refinements in our understanding of the observational properties of these objects, including their dust properties (Brooks et al., 2024; Setton et al., 2025; Xiao et al., 2025; Ferrara, 2024; Ferrara et al., 2025) and Balmer absorption and emission (Setton et al., 2024; Naidu et al., 2025; D’Eugenio et al., 2025; de Graaff et al., 2025), there is still a high degree of uncertainty about their nature. This is due to the complex interplay between a variety of their physical properties and the multitude of observational constraints that have been identified for these objects.
@@
Future work in this area should examine the impacts of both the computational constraints and model choices made in simulation. Techniques like zoom-in simulations could be used to create mock observations of potential dim LRDs at different resolutions to determine if their absence in Astrid is the result of resolution effects. If such simulations can reduce the range of resolution effects well below the proposed size of star-dominated LRDs (100-200 pc) they would also provide an ideal test bed for evaluating the viability of different models for such LRDs. Similarly, simulations run with different black hole models could provide insight into the conditions necessary for AGN-dominated or hybrid LRD formation. For example, mock observations of galaxies in different simulations within the BRAHMA simulation suite (Bhowmick et al., 2024), which feature different black hole seeding models, would provide insights into the potential impacts of over-massive black holes on the properties and number density of LRDs. The findings of these works would also help inform the spatial resolution and black hole seeding models used in the next generation of cosmological simulations.

### v2

@@ 5 Where Are The Dimmer LRDs?
The population of LRDs in Astrid represents a subset of the observed population, as their number density is significantly lower than those found in observations, and they are concentrated within the brighter portion of the observed LRD population. Astrid contains $\sim 1/100$ of the expected number of LRDs with UV magnitudes brighter than $\rm M_{UV}=-19$, but completely lacks LRDs with $\rm M_{UV}\geq-18.5$ resulting in an overall LRD number density of $\sim 1/1000$ of observations (Kocevski_2024; Kokorev_2024; Perez_Gonzalez_2024).
Figure 11: Spectrum of a $M_{*}=10^{9.4}\,M_{\odot}$ galaxy hosting a black hole with mass $M_{BH}=10^{6.1}\,M_{\odot}$, accreting at 116% the Eddington rate, with the dust occlusion on each star increased by a factor of four and applying an AGN SED model with a Hydrogen column density of $N_{H}=10^{24}$ $\rm N/cm^{2}$. This is the same galaxy that appears in the bottom left panel of Figure 6. For comparison we include the “Total SED” from that figure (with the original AGN SED and stellar dust models) as a purple line on this plot. With these changes, the galaxy fits the criteria to be an LRD.
@@
This galaxy has an F277W-F [^4] W color of 0.88, which is insufficient to host an LRD. The primary differences between the spectrum of this galaxy and that of the example LRD in Figure 2 are the relative brightness of the AGN and the dust attenuation of both components. In the LRD spectrum, before dust attenuation, the AGN and stellar components have similar brightness in the F [^4] W band. Despite having one of the brightest AGN relative to its stellar mass among the lower stellar mass galaxies, the AGN emission of the dim LRD candidate galaxy is only $\sim 16\%$ as bright as its host galaxy in the F [^4] W band before dust attenuation. This indicates very few, if any, of the lower mass galaxies in Astrid have AGN that are bright enough to significantly impact their overall color. This is likely the result of both black hole accretion and star formation being induced by the presence of cold dense gas, producing host galaxies which are brighter due to their increased stellar mass, and young stellar population. This coupling is broken once AGN feedback begins quenching the host galaxy, as we see in our LRDs and models of high redshift galaxies hosting over-massive black holes (Pacucci_Loeb_2024; Pacucci_2024).
@@
These two factors are heavily impacted by modeling choices in our mock observation pipeline. AGN bolometric luminosity is determined by the simulation, but the AGN SED can impact how that luminosity is distributed, and is dependent on the properties of the gas very near the black hole, which is not resolved by cosmological simulations. Recent work (e.g., Naidu_2025), suggests that AGN embedded in a denser gas cloud, similar to models of direct collapse black holes (e.g. Pacucci_2015; Pacucci_2016), may be a strong candidate for LRD emission, which produces a significantly different SED than our more traditional model. Similarly, our dust model has a normalization factor calibrated in astrid_galaxy_formation, but may not be tuned appropriately for LRD galaxies.
@@
We experiment with these two modeling choices on our dim LRD candidate galaxy. We increase the hydrogen column density of our AGN SED model in cloudy from $10^{23}$ $\rm N/cm^{2}$ to $10^{24}$ $\rm N/cm^{2}$ to approximate the effect of the ”gas-enshrouded" AGN models proposed in works like Inayoshi_2025; Naidu_2025 and, Taylor_2025. Additionally, we increase the dust normalization by a factor of four (from $\kappa_{\rm ISM}=10^{4.1}$ to $\kappa_{\rm ISM}=10^{4.7}$) for the stellar emission to explore the impact of errors in normalization. This also serves as an approximate test of the impact of potential under-density due to simulation effects like resolution limitations and gravitational softening.
The spectrum of the dim LRD candidate galaxy with these changes is shown in Figure 11. This galaxy does appear as an LRD with these changes, as its $\rm F277W-F444W$ color has shifted from 0.88 to 1.16. With an F [^4] W magnitude of 25.95, and a stellar mass of $\rm M_{*}=10^{9.41}\,M_{\odot}$ it would be the dimmest and lowest mass LRD in our population. This indicates it is possible for lower mass galaxies hosting Eddington accreting black holes in simulations to appear as LRDs with specific modeling choices.
The “gas-enshrouded" AGN SED introduces another reddening mechanism to these sources, which could also address the lack of very red LRDs in Astrid. Given the “gas-enshrouded" AGN model we adopted in this experiment was fairly conservative (Naidu_2025 found a Hydrogen column density of $\rm N(H)/cm^{2}\sim 10^{26}$) it is likely that a more extreme version of this model would have a significantly redder spectrum. If such an SED were applied to Eddington or Super-Eddington sources, they may appear similar to the reddest objects seen in observations. We intend to explore these alternative AGN models in more depth in future work.
@@
The increased hydrogen column density in the AGN SED introduces both a steeper intrinsic AGN SED and a significant Balmer break, both of which contribute to the red appearance of this galaxy. These qualities are especially important for any LRDs with minimal dust content, like those suggested by Setton_2025; Xiao_2025, and would result in the consistent Balmer break transition that is found in Setton_2024.
@@
The increased stellar dust attenuation has two primary effects: (I) it reddens the stellar spectrum and (II) it decreases the brightness of the stellar component, which increases the relative contribution of the AGN to the total emission of the source. This increase in AGN contribution is particularly impactful given the significant red color of the “gas-enshrouded" AGN SED. If these high Eddington black holes were present in dimmer, lower mass host galaxies, that would also produce sources with heightened AGN contributions. With the black hole models used in Astrid this is unlikely to occur due to the black hole seeding and feedback models, but it may be possible under alternative models. The BRAHMA simulation suite (Bhowmick_2024) suggests that the over-massive black holes seen in observations are reproducible in simulation with specific seeding techniques, and that much of their mass is the result of black hole mergers. This would allow for the formation of more massive black holes in lower mass galaxies, as their mass growth would be less coupled to star formation. These more massive black holes would be capable of brighter AGN emission without exceeding the Eddington limit, potentially appearing as LRDs.
@@ 6.1 Discussion
The size of the Astrid simulation is exceptionally advantageous for works such as this one, as it provides a cosmological volume that is much larger than most observational datasets. Given the density of LRDs found by observations across the redshifts we analyze in this work, we would expect to identify tens of thousands of LRDs in our dataset (Kocevski_2024; Kokorev_2024; Matthee_2023; Perez_Gonzalez_2024). Despite this, we find only 17 LRDs, and 24 no-break LRDs in our dataset, and those galaxies would be among the brightest LRDs observed.
Due to the compact nature of LRDs, the computational constraints that arise on small scales in cosmological simulations may limit their ability to form. The SPH method used in Astrid allows for highly accurate physical modeling on a wide range of scales. Still, it is limited at the smallest scales, due to the mass resolution of particles and gravitational softening. This is relevant for all LRD models, but is of particular importance to the star-dominated models because they are reliant on the existence of extremely dense star-forming regions, which can push the theoretical limits of the current cosmological model (Guia_2024; Ma_2024; Leung_2024; Akins_2024).
Simulations also must choose specific physical models for processes regarding black holes (e.g., seeding, accretion, and merging) and gas dynamics/radiative properties (e.g., AGN feedback, star formation), etc., all of which impact the possibility of an LRD forming. The choices for these models in Astrid are described in detail in astrid_BHs; astrid_galaxy_formation and have been shown to have results consistent with observations and other simulations on a wide variety of metrics. AGN-dominated fits of observed LRDs often suggest that over-massive black holes must be present to produce AGNs with sufficient brightness (Pacucci_Loeb_2024; Pacucci_2023; Inayoshi_2024; Durodola_2024), but most current black hole seeding and accretion models, including those used in Astrid  do not have the pathways to create such over-massive black holes.
@@
Despite refinements in our understanding of the observational properties of these objects, including their dust properties (Brooks_2024; Setton_2025; Xiao_2025; Ferrara_2024; Ferrara_2025_a) and Balmer absorption and emission (Setton_2024; Naidu_2025; DEugenio_2025; deGraff_2025), there is still a high degree of uncertainty about their nature. This is due to the complex interplay between a variety of their physical properties and the multitude of observational constraints that have been identified for these objects.
@@
Future work in this area should examine the impacts of both the computational constraints and model choices made in simulation. Techniques like zoom-in simulations could be used to create mock observations of potential dim LRDs at different resolutions to determine if their absence in Astrid is the result of resolution effects. If such simulations can reduce the range of resolution effects well below the proposed size of star-dominated LRDs (100-200 pc) they would also provide an ideal test bed for evaluating the viability of different models for such LRDs. Similarly, simulations run with different black hole models could provide insight into the conditions necessary for AGN-dominated or hybrid LRD formation. For example, mock observations of galaxies in different simulations within the BRAHMA simulation suite (Bhowmick_2024), which feature different black hole seeding models, would provide insights into the potential impacts of over-massive black holes on the properties and number density of LRDs. The findings of these works would also help inform the spatial resolution and black hole seeding models used in the next generation of cosmological simulations.